OpenStudy (anonymous):

When there are 30 girls in a dance, there are 12 boys in the dance. When there are 45 girls in the dance, there are 17 boys in the dance. Assume the relationship is linear. Which equation models the relationship between the number of girls, x, and the number of boys, y? y = 3x + 26 y=1/2x+22 y=1/3x+2 y=3x+9

OpenStudy (danjs):

x = #girls y = # boys they give you two points on the line (x,y) = (30,12) (x1,y1) = ( 45,17) calculate the slope from those 2 points first

OpenStudy (danjs):

Then use one of the points in the slope-point formula for a line y - y1 = m*(x-x1) to get your equation for the line

OpenStudy (anonymous):

oh i was doing it all wrong i was plugging those in for x and y

OpenStudy (danjs):

nah, those are each a point on the line

OpenStudy (anonymous):

whats the formula for finding the slope?

OpenStudy (danjs):

Given 2 points above (x,y) and (x1,y1) \[slope =m=\frac{ \Delta y }{ \Delta x }= \frac{ y-y1 }{ x-x1 }\]

OpenStudy (anonymous):

ok so i plug those numbers into this formula?

OpenStudy (danjs):

Right, it is just the difference in the y coordinate , divided by, the difference in the x coordinate

OpenStudy (anonymous):

ok im gonna try to solve it

OpenStudy (anonymous):

how exactly do i plot for y1 like do i put the number and 1?

OpenStudy (danjs):

\[y1 = y _{1}\] just a different y value, was lazy to write the subscript

OpenStudy (anonymous):

oooooh ok its ok thanks

OpenStudy (danjs):

What are you getting as the slope?

OpenStudy (anonymous):

great work guys u can be the best in ur class

OpenStudy (anonymous):

thank you noland

OpenStudy (danjs):

(x,y) = (30,12) (x1,y1) = ( 45,17) you getting the slope? m = (17 - 12) / (45 - 30) =

OpenStudy (anonymous):

yes im working on the problem right now and i was doing it wrong

OpenStudy (danjs):

what did you do?

OpenStudy (danjs):

you will get teh same slope if you do it the other way, m = (12 - 17) / ( 30 - 45)

OpenStudy (anonymous):

how r u guys doing so far?

OpenStudy (anonymous):

i did 12/30 =12-12(1) 30-30(1)

OpenStudy (danjs):

(x,y) = (30,12) (x1,y1) = ( 45,17) \[m = \frac{ y-y1 }{ x-x1 } = \frac{ 12 - 17 }{ 30-45 } = \frac{ -5 }{ -15 } = \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

for (12-17) i got -5 and for (30-45) i got -15

OpenStudy (danjs):

yep, then reduce that fraction like above, to 1/3 by dividing both top and bottom by -5

OpenStudy (danjs):

Now that you have the slope; m = 1/3 and point on the line (x1,y1) = ( 45,17) Use y - y1 = m*(x-x1) to form an equation for the line

OpenStudy (anonymous):

ok why do we divide -5 into it?

OpenStudy (danjs):

reduce the fraction -5/-15 = 1 / 3

OpenStudy (anonymous):

oh ok

OpenStudy (danjs):

now just sub into the point-slope form y - y1 = m*(x-x1) using slope m = 1/3 and any of the two points

OpenStudy (danjs):

m = 1/3 (x1,y1) = (45,17) y - y1 = m*(x-x1) y - 17 = (1/3)*(x - 45)

OpenStudy (anonymous):

ok im getting why i cant do this im not plugging them in right sorry if this may be fustrating for you

OpenStudy (danjs):

now just have to solve that for y = .... y - 17 = (1/3)*(x - 45) y - 17 = 1/3 x - 15 y = 1/3 x + 2

OpenStudy (danjs):

nah, you are learning

OpenStudy (anonymous):

thank you alot!

OpenStudy (danjs):

if you have another example you want to go through, you can tag me in it

OpenStudy (anonymous):

ok i will if i find one i fanned you aswell

OpenStudy (danjs):

i would go back through this thread, and write on paper what i typed in, it may make more sense

OpenStudy (anonymous):

ok i will i think that will help me if i ever see a problem like this again

OpenStudy (danjs):

yep, just have to realize the problem is giving you 2 points, from there you just find the slope, and use point-slope form to make a line

OpenStudy (anonymous):

ok thanks again truely grateful